OpenStudy (appleduardo):

which is the limit of the following function? could somebody help me please?

OpenStudy (appleduardo):

\[Lim \frac{ 6 }{x- 5 } \]

OpenStudy (appleduardo):

when x-->5

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5}\frac{6}{x-5}\] the answer is infinity. 6/0, zero goes into six infinitely many times.

OpenStudy (appleduardo):

so how can i express the result? i mean which is the limit then? is jut 6/0 ?

OpenStudy (anonymous):

you would just write the equation equals infinity (the sideways 8) \[lim_{x \rightarrow 5}\frac{6}{x-5}=\infinity\]

OpenStudy (appleduardo):

oh! thank you so much! :P